Laurel Electronics reported the following information at its annual meeting: The company had cash and marketable securities worth $1,235,455, accounts payables worth $4,159,357, inventory of $7,136,200, accounts receivables of $3,450,200, short-term notes payable worth $1,146,500, and other current assets of $121,455. What is the company's net working capital?
Answer #1

Total current assets=cash and marketable securities+inventory+AR+other current assets

=(1,235,455+7,136,200+3,450,200+121,455)=$11943310

Total current liabilities=AP+Short term notes payable

=4,159,357+1,146,500=$5305857

Net working capital=Total current assets-Total current liabilities

=11943310-5305857

=$6637453